OpenStudy (anonymous):

Can someone help me with this problem solve equation -3(m-4)+2(5+2m)=29

OpenStudy (anonymous):

what grade is it?

OpenStudy (anonymous):

College math 115

OpenStudy (anonymous):

Okay, Use distribute property(mulitply out)

OpenStudy (anonymous):

Thanks ! :). But I don't know how to solve it. I tried and keep getting the wrong answer.

OpenStudy (anonymous):

I will do the first term -3m+12

OpenStudy (anonymous):

you try other terms

OpenStudy (anonymous):

I got -3m+12+ 10+4m=29. Then I get lost lol

OpenStudy (anonymous):

you are right so far

OpenStudy (anonymous):

now add like terms

OpenStudy (anonymous):

Ok thanx for your help! :)

OpenStudy (anonymous):

If I helped you somehow, fan me

OpenStudy (anonymous):

Yea. It's m=7. *-)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Can you help me with just one more this one is a for sure challenge

OpenStudy (anonymous):

7m-(2m-9)=39

OpenStudy (anonymous):

is it 5m-9=39?